# Modeling Time-Variant Responses of Optical Compressors with Selective State Space Models

### v2

@@ 5.1 Unseen settings and sampling density
Table 7: Loss and performance metrics for models trained on Softube CL 1B time datasets with different parameter sampling densities. The ’Set’ columns indicate the training sets used, corresponding to the suffix of the training set names detailed in Table 3. The ’EP’ columns reports the last training epoch at which the models exhibited improvements. Models with the lowest loss are highlighted in bold, and the minimum value of each metric is in italics.
Set Model MSE MAE ESR RMSE SFE M-STFTE PSM EP
0.20 S6 $\mathit{3.32\cdot 10^{-4}}$ $\mathit{8.83\cdot 10^{-3}}$ $\mathit{2.43\cdot 10^{-1}}$ $8.48\cdot 10^{-3}$ $\mathit{7.18\cdot 10^{-3}}$ $2.00$ $0.94$ 60
S4D $3.49\cdot 10^{-4}$ $8.85\cdot 10^{-3}$ $2.55\cdot 10^{-1}$ $8.20\cdot 10^{-3}$ $7.73\cdot 10^{-3}$ $\mathit{1.93}$ $0.96$ 53
ED $3.52\cdot 10^{-4}$ $8.36\cdot 10^{-3}$ $2.57\cdot 10^{-1}$ $\mathit{7.94\cdot 10^{-3}}$ $8.63\cdot 10^{-3}$ $1.95$ $0.95$ 43
LSTM $6.83\cdot 10^{-4}$ $1.27\cdot 10^{-2}$ $5.31\cdot 10^{-1}$ $1.23\cdot 10^{-2}$ $8.53\cdot 10^{-3}$ $3.31$ $0.95$ 6
0.10 S6 $\mathit{4.16\cdot 10^{-4}}$ $\mathit{9.27\cdot 10^{-3}}$ $\mathit{2.65\cdot 10^{-1}}$ $8.90\cdot 10^{-3}$ $9.38\cdot 10^{-3}$ $2.11$ $0.94$ 39
S4D $5.78\cdot 10^{-4}$ $1.07\cdot 10^{-2}$ $3.68\cdot 10^{-1}$ $1.01\cdot 10^{-2}$ $\mathit{8.73\cdot 10^{-3}}$ $2.39$ $0.96$ 11
ED $4.32\cdot 10^{-4}$ $9.15\cdot 10^{-2}$ $2.75\cdot 10^{-1}$ $8.78\cdot 10^{-3}$ $1.04\cdot 10^{-2}$ $2.14$ $0.95$ 19
LSTM $4.37\cdot 10^{-4}$ $9.14\cdot 10^{-3}$ $2.78\cdot 10^{-1}$ $\mathit{8.61\cdot 10^{-3}}$ $9.42\cdot 10^{-3}$ $\mathit{2.01}$ $0.95$ 12
0.05 S6 $5.27\cdot 10^{-4}$ $1.10\cdot 10^{-2}$ $3.86\cdot 10^{-1}$ $1.07\cdot 10^{-2}$ $1.20\cdot 10^{-2}$ $2.62$ $0.92$ 4
S4D $\mathit{3.71\cdot 10^{-4}}$ $8.87\cdot 10^{-3}$ $\mathit{2.71\cdot 10^{-1}}$ $8.40\cdot 10^{-3}$ $7.98\cdot 10^{-3}$ $\mathit{2.01}$ $0.95$ 33
ED $4.33\cdot 10^{-4}$ $\mathit{8.78\cdot 10^{-3}}$ $3.17\cdot 10^{-1}$ $\mathit{8.31\cdot 10^{-3}}$ $9.22\cdot 10^{-3}$ $2.11$ $0.95$ 3
LSTM $3.99\cdot 10^{-4}$ $8.97\cdot 10^{-3}$ $2.92\cdot 10^{-1}$ $8.44\cdot 10^{-3}$ $\mathit{7.71\cdot 10^{-3}}$ $2.07$ $0.95$ 11
@@ 5.2 Time behavior
Results indicate that the combination of fast attack and slow release is the most challenging compression scenario to emulate, yielding the highest errors across all proposed models trained on individual compression settings. On the other hand, slow attack scenarios pose significant challenges for the baseline models. In addition, the performance degrades markedly when comparing models trained exclusively on settings including slow attack or slow release to models trained on the four combinations. As observed in previous experiments, these outcomes suggest that learning variable attack times across the wide range of CL 1B is especially difficult. While all models show varying degrees of increased and decreased errors depending on the conditioning, they generally learn the conditioning cases to different extents, and the learning process seems to average out accuracy among them. Notably, S6 and LSTM models conditioned to handle multiple compression timing behaviors do not always perform worse; instead, they distribute errors across the examples, performing poorer in some settings and better in others. In particular, they improve the accuracy when both parameters are set to slow times or both are set to fast ones but decrease for the other conditions. The ED and S4D variations of the proposed architecture exhibit greater susceptibility to degradation, resulting in higher errors across all settings. In comparison, the baseline models generally demonstrate inferior performance, presenting more difficulties for slow attack settings than for fast ones. The proposed architecture with the S6 block consistently delivers the best performance, whereas the TCN-b model performs notably poorly, indicating a potential difficulty in learning varied timing behaviors and a tendency to favor for settings with rapid attack and release times. The low error rates might also be influenced by the limited dataset used for training, given the model’s reliance on convolutional layers.
@@ 6 CONCLUSION
The most challenging scenarios involved variable temporal behavior, where datasets of smaller size provided the best results. For the target devices used in this work, the exact parameter values are often not known; they are neither marked around the knobs nor reported in the manuals. Only the extreme values and some intermediate points are typically indicated. We also lack precise knowledge about whether parameter changes occur on a linear or nonlinear scale. Therefore, we opted to sample the parameters using equally spaced intervals, although employing a more sophisticated sampling strategy could potentially yield better outcomes. When evaluating the models’ performance on individual compression settings, the combination of fast attack with slow release proved to be the most challenging to emulate. The complexity arises from allowing a very brief initial segment of the sound to pass unaltered before gain reduction begins, which may lead to abrupt dynamic changes within a short time frame. Conversely, when training the models to emulate various compression timing behaviors, there is an expected overall decrease in accuracy. This general reduction in precision is particularly noticeable for settings with slow attack and release times.
Finally, in the conditioning block, we utilize information from the signal’s frequency domain based on Fourier analysis. This approach aids in capturing the global properties of a series of input signal segments, providing detailed information about the input signal’s energy, which influences the compression process. In this regard, alternative frequency domain representations, such as wavelet transforms, could provide a further advantage. They might provide a more effective representation encompassing the frequency and time domains, enabling efficient access to the signal’s localized information and potentially enhancing the model’s accuracy.

### v3

@@ 5.1 Unseen settings and sampling density
Table 8: Loss and performance metrics for models trained on Softube CL 1B time datasets with different parameter sampling densities. The ‘Set’ columns indicate the training sets used, corresponding to the suffix of the training set names detailed in Table 3. Models with the lowest loss and the minimum value for each metric are highlighted in bold.
Set Model MSE MAE ESR RMSE SFE M-STFTE PSM
0.20 S6 $\mathbf{3.32\cdot 10^{-4}}$ $\mathbf{8.83\cdot 10^{-3}}$ $\mathbf{2.43\cdot 10^{-1}}$ $8.48\cdot 10^{-3}$ $\mathbf{7.18\cdot 10^{-3}}$ $2.00$ $0.94$
S4D $3.49\cdot 10^{-4}$ $8.85\cdot 10^{-3}$ $2.55\cdot 10^{-1}$ $8.20\cdot 10^{-3}$ $7.73\cdot 10^{-3}$ $\mathbf{1.93}$ $0.96$
ED $3.52\cdot 10^{-4}$ $8.36\cdot 10^{-3}$ $2.57\cdot 10^{-1}$ $\mathbf{7.94\cdot 10^{-3}}$ $8.63\cdot 10^{-3}$ $1.95$ $0.95$
LSTM $6.83\cdot 10^{-4}$ $1.27\cdot 10^{-2}$ $5.31\cdot 10^{-1}$ $1.23\cdot 10^{-2}$ $8.53\cdot 10^{-3}$ $3.31$ $0.95$
0.10 S6 $\mathbf{4.16\cdot 10^{-4}}$ $\mathbf{9.27\cdot 10^{-3}}$ $\mathbf{2.65\cdot 10^{-1}}$ $8.90\cdot 10^{-3}$ $9.38\cdot 10^{-3}$ $2.11$ $0.94$
S4D $5.78\cdot 10^{-4}$ $1.07\cdot 10^{-2}$ $3.68\cdot 10^{-1}$ $1.01\cdot 10^{-2}$ $\mathbf{8.73\cdot 10^{-3}}$ $2.39$ $0.96$
ED $4.32\cdot 10^{-4}$ $9.15\cdot 10^{-2}$ $2.75\cdot 10^{-1}$ $8.78\cdot 10^{-3}$ $1.04\cdot 10^{-2}$ $2.14$ $0.95$
LSTM $4.37\cdot 10^{-4}$ $9.14\cdot 10^{-3}$ $2.78\cdot 10^{-1}$ $\mathbf{8.61\cdot 10^{-3}}$ $9.42\cdot 10^{-3}$ $\mathbf{2.01}$ $0.95$
0.05 S6 $5.27\cdot 10^{-4}$ $1.10\cdot 10^{-2}$ $3.86\cdot 10^{-1}$ $1.07\cdot 10^{-2}$ $1.20\cdot 10^{-2}$ $2.62$ $0.92$
S4D $\mathbf{3.71\cdot 10^{-4}}$ $8.87\cdot 10^{-3}$ $\mathbf{2.71\cdot 10^{-1}}$ $8.40\cdot 10^{-3}$ $7.98\cdot 10^{-3}$ $\mathbf{2.01}$ $0.95$
ED $4.33\cdot 10^{-4}$ $\mathbf{8.78\cdot 10^{-3}}$ $3.17\cdot 10^{-1}$ $\mathbf{8.31\cdot 10^{-3}}$ $9.22\cdot 10^{-3}$ $2.11$ $0.95$
LSTM $3.99\cdot 10^{-4}$ $8.97\cdot 10^{-3}$ $2.92\cdot 10^{-1}$ $8.44\cdot 10^{-3}$ $\mathbf{7.71\cdot 10^{-3}}$ $2.07$ $0.95$
@@ 5.2 Time behavior
Results indicate that the combination of fast attack and slow release is the most challenging compression scenario to emulate, yielding the highest errors across all proposed models trained on individual compression settings. On the other hand, slow attack scenarios pose significant challenges for the baseline models. In addition, the performance degrades markedly when comparing models trained exclusively on settings including slow attack or slow release to models trained on the four combinations. As observed in previous experiments, these outcomes suggest that learning variable attack times across the wide range of CL 1B is especially difficult.
While all models show varying degrees of increased and decreased errors depending on the conditioning, they generally learn the conditioning cases to different extents, and the learning process seems to average out accuracy among them. Notably, S6 and LSTM models conditioned to handle multiple compression timing behaviors do not always perform worse; instead, they distribute errors across the examples, performing poorer in some settings and better in others. In particular, they improve the accuracy when both parameters are set to slow times or both are set to fast ones but decrease for the other conditions.
The ED and S4D variations of the proposed architecture exhibit greater susceptibility to degradation, resulting in higher errors across all settings. In comparison, the baseline models generally demonstrate inferior performance, presenting more difficulties for slow attack settings than for fast ones. The proposed architecture with the S6 block consistently delivers the best performance, whereas the TCN-b model performs notably poorly, indicating a potential difficulty in learning varied timing behaviors and a tendency to favor for settings with rapid attack and release times. The low error rates might also be influenced by the limited dataset used for training, given the model’s reliance on convolutional layers.
@@ 6 CONCLUSION
The most challenging scenarios involved variable temporal behavior, where datasets of smaller size provided the best results. For the target devices used in this work, the exact parameter values are often not known; they are neither marked around the knobs nor reported in the manuals. Only the extreme values and some intermediate points are typically indicated. We also lack precise knowledge about whether parameter changes occur on a linear or nonlinear scale. Therefore, we opted to sample the parameters using equally spaced intervals, although a more sophisticated sampling strategy could potentially yield better outcomes.
When evaluating the models’ performance on individual compression settings, the combination of fast attack with slow release proved to be the most challenging to emulate. The complexity arises from allowing a very brief initial segment of the sound to pass unaltered before gain reduction begins, which may lead to abrupt dynamic changes within a short time frame. Conversely, when training the models to emulate various compression timing behaviors, there is an expected overall decrease in accuracy. This general reduction in precision is particularly noticeable for settings with slow attack and release times.
Finally, in the conditioning block, we utilize information from the signal’s frequency domain based on Fourier analysis. This approach aids in capturing the global properties of a series of input segments, providing information about the input’s energy, which influences the compression process. In this regard, alternative frequency domain representations, such as wavelet transforms, could provide a further advantage. They might provide a more effective representation encompassing the frequency and time domains, enabling efficient access to the signal’s localized information and potentially enhancing the model’s accuracy.
Further improvements in the model’s accuracy could be achieved by designing a loss function that places greater emphasis on errors during sharp input transients, which are particularly challenging to model accurately, especially in the presence of fast attack times. The current loss function, which averages errors across extended temporal segments of the output signal, may be insufficiently sensitive to such errors. Although reducing the training batch size could help address this issue, it would likely lead to a significant increase in training time.